when 286g CU reacts with an excess of HNO3 what mass of NO is f


1
Expert's answer
2020-11-09T14:00:15-0500

"\begin{aligned}
\small 3Cu_{(s)} + 8HNO_{3(aq)} \to 2NO_{(g)} + 3Cu(NO_3)_{2(aq)} + 4H_2O_{(l)}
\end{aligned}"

from the reaction above,

3 moles of Cu produces 2 moles of NO.


286g of Cu = 4.5 moles of Cu.

4.5 moles of Cu produces "(\frac{4.5×2}{3}=)" 3 moles of NO



3 moles of NO = 3(14+16) = 3(30) = 90g.


"\therefore" when 286g CU reacts with an excess of HNO3, 90g mass of NO is produced.